OpenStudy (anonymous):

Suppose that f has an inverse and f(−4)=3, f′(−4)=4/7. If g=(1)/(f−1), what is g′(3)

OpenStudy (freckles):

what does 1f-1 mean?

OpenStudy (freckles):

i wonder if you just mean that g is the inverse of f ?

OpenStudy (freckles):

\[\text{ if so } g'(3)=\frac{1}{f'(g(3))}\]

OpenStudy (freckles):

and if f(-4)=3 then g(3)=-4

OpenStudy (freckles):

if f and g are inverses

OpenStudy (anonymous):

g=(1)/(f−1)

OpenStudy (freckles):

\[g(x)=\frac{1}{f(x)-1}?\]

OpenStudy (freckles):

Use quotient rule